find the slope and the y-intercept of the line. 5x=3y+15

Found 2 solutions by mananth, cira554:
Answer by mananth(16946)

5x=3y+15
5x-15=3y
/3
(5/3)x-5=y
y=(5/3)x-5
m= slope = 5/3
y intercept =-5

Answer by cira554(7)

5x=3y+15 rearranging it we get 3y=5x-15,
dividing by 3 all through we get y=5/3x-5 which is in the form y=mx+c,
m is our slope and c our y-intercept.
therefore from our equation the slope is 5/3 and the y-intrcept=-5